(a) In one cycle of a heat engine: 70 J are input by working, 90 J are output by working, 100 J are input by heating, 80 J are output by heating. Write down which one of the following is true: (A) İWnet 1-160 J (B) IWnet-20 J (C) η= 0.25 (D) IWnet--20 J (E) η =-20%
Answer #1

Work net=work output-work input=90-70=20 Joule.

Heat supplied or Heat input =100.

Efficiency= Work done / Heat supplied=20/100=.20

Therefore efficiency is 20%.

Option B is correct .